For a binomial process, the probability of success is 40 percent and the number of trials is 5. Find P(X ≤ 1).
Answer #1

P(p) = 0.4

n =5

P(X<=1) = P(X=0)+P(X=1)

= 5C0 *(0.4^5) + 5C1* (0.4^4)*(0.6) = 0.337